OpenStudy (anonymous):

Logarithm question:

OpenStudy (goformit100):

????

OpenStudy (anonymous):

\[\log _{9}(9^6/27)\]

OpenStudy (anonymous):

=\[\log_{9}(9^{6}/27)\] \[\log_{3^{2}}(3^{12}/3^{3})\] \[\log_{3^{2}}(3^{12-3})\] \[9/2log_{3}(3)\] \[9/2*1\] = 9/2

OpenStudy (anonymous):

thank you :)